chickens and rabbits were placed under the same tent. The chickens and the rabbits have a total of 360 legs. If there were 105 animals, how many were rabbits?
Answer by Earlsdon(6287)

Let C = the number of chickens and R = the number of rabbits.
Then: 2C + 4R = 360 The total number of legs.
and: C + R = 105 The total number of animals. Rewrite this as: C = 105 - R and substitute into the first equation.
2(105 - R) + 4R = 360 Simplify and solve for R, the number of rabbits.
210 - 2R + 4R = 360 Add the R's.
210 + 2R = 360 Subtract 210 from both sides.
2R = 150 Divide both sides by 2.
R = 75 There were 75 rabbits.